The length of a rectangular garden is 3meter greater than the width. If the area is 88m2, what is its dimension?
Answer by mananth(16949)

let width be x
length = x+3
Area of rectangle = L*W
x*(x+3)=88
x^2+3x=88
x^2+3x-88 =0
x^2+11x-8x-88=0
x(x+11)-8(x+11)=0
(x+11)(x-8)=0
x=-11 OR x=8
ignore negative
x=8 meters
length = 8+3=11 meters